OpenStudy (anonymous):

what are the zeros of the polynomial function f(x)=(x+6)(x+8)(x+15)

jimthompson5910 (jim_thompson5910):

f(x)=(x+6)(x+8)(x+15) 0=(x+6)(x+8)(x+15) (x+6)(x+8)(x+15) = 0 now use the zero product property